# Physics-aware Machine Learning Revolutionizes Scientific Paradigm for Machine Learning and Process-based Hydrology

### v4

@@ 2.2.1 Physics-informed Neural Networks
PDE solutions are represented as neural networks by including the square of the PDE residual in the loss function, resulting in a neural network-based PDE solver [24]. Recently, this approach is refined further and called “physics-informed neural networks (PINNs) [25],” initiating a flurry of follow-up work. For example, we consider a parametrized PDE system given by
$\begin{gathered}f\left(\mathbf{x},t,u,\frac{\partial u}{\partial\mathbf{x}},\frac{\partial^{2}u}{\partial\mathbf{x^{2}}},\frac{\partial u}{\partial t},\ldots,\vartheta\right)=0,\quad\mathbf{x}\in\Omega,t\in[0,T]\\ u\left(\mathbf{x},t_{0}\right)=g_{0}(\mathbf{x})\quad\mathbf{x}\in\Omega\quad u(\mathbf{x},t)=g_{\Gamma}(t),\quad\mathbf{x}\in\partial\Omega,t\in[0,T],\end{gathered}$ (1)
where $u:[0,T]\times\mathbb{X}\rightarrow\mathbb{R}^{n}$ is the solution, subject to initial condition $g_{0}(\mathbf{x})$ and boundary condition $g_{\Gamma}(t)$, which can be of various types such as periodic, Dirichlet, or Neumann. The PDE parameters are denoted by $\vartheta$ = [$\vartheta 1$, $\vartheta 2$, . . . ]. The residual of the PDE is represented by $f$, which includes the differential operators (i.e., $\frac{\partial u}{\partial\mathbf{x}},\frac{\partial^{2}u}{\partial\mathbf{x^{2}}},\frac{\partial u}{\partial t}$, . . . ). The physical domain is denoted by $\Omega$, with its corresponding boundary represented by $\partial\Omega$.
@@
$\begin{gathered}\min_{\theta}\mathcal{L}_{\mathrm{PINN}}(\theta)=w_{i}\mathcal{L}_{IC}+w_{b}\mathcal{L}_{BC}+w_{d}\mathcal{L}_{\text{Data }}+w_{p}\mathcal{L}_{PDE}\\ \mathcal{L}_{IC}=\|u\left(\mathbf{x},t_{0}\right)-g_{0}(\mathbf{x})\|_{\Omega}\\ \mathcal{L}_{BC}=\|u(\mathbf{x},t)-g_{\Gamma}(t)\|_{\partial\Omega}\\ \mathcal{L}_{\text{Data }}=u|_{\Omega}-\hat{u}|_{Data}\\ \mathcal{L}_{PDE}=\|f\left(\mathbf{x},t,u,\frac{\partial u}{\partial\mathbf{x}},\frac{\partial^{2}u}{\partial\mathbf{x^{2}}},\frac{\partial u}{\partial t},\ldots,\vartheta\right)\|_{\Omega},\end{gathered}$ (2)
@@
In current ML-based methods for scientific computing, vanilla Physics-Informed Neural Networks (PINNs) are typically formulated in a pointwise manner, as depicted in Fig. 3(a). These methods utilize neural networks as the ansatz for the solution function and optimize a loss function to minimize the deviation from the given equations, leveraging auto-differentiation for exact, mesh-free derivatives. One of the primary advantages of vanilla PINNs is their ability to avoid the need for a discretization grid, making them particularly useful for solving inverse and higher-dimensional problems. However, despite these advantages, vanilla PINNs exhibit lower efficiency compared to classical methods in solving most Partial Differential Equations (PDEs), for several reasons [146, 147, 148, 149]: (1) the challenging optimizations due to soft physics or PDE constraints; (2) the difficulty of conveying information from the initial or boundary conditions to unseen interior regions or to future times; (3) the sensitivity to hyperparameter selection; (4) the slow training/re-optimization process in practical applications and different instances; and (5) failure to converge on time-dependent PDEs or high-frequency, multi-scale PDEs, as shown in Fig. 3(b).
As shown in Table 2.2, a vast and growing body of work aims to address these challenges. First, many problem-specific insights are utilized by improving the training of PINNs. For example, Universal PINNs [104] improve the PINNs by a lower bound constrained uncertainty weighting algorithm and a multi-scale deep neural network. A distinct conservative physics-informed neural network (cPINN) [105] is devised for discrete domains dealing with nonlinear conservation laws. Another approach, nPINNs [106], extends PINNs to handle parameter and function inference for integral equations like nonlocal Poisson and nonlocal turbulence models. An improved PINN approach [107] is presented to solve localized wave solutions of the derivative nonlinear Schrödinger equation in complex space, achieving faster convergence and optimal simulation performance. An automatic numerical solver [108] is proposed for the Allen-Cahn and Cahn-Hilliard equations employing advanced PINNs. A straightforward and efficient characteristic-informed neural network [109] is developed for solving forward and inverse problems in hyperbolic PDEs. NSFnets [110] is proposed for the incompressible Navier-Stokes equations using automatic differentiation. Furthermore, for solving Navier-Stokes equations on irregular geometries, a geometry-aware physics-informed neural network (GAPINN) [111] is presented. In addition, a coupled physics-informed neural networks (CPINNs) [112] is designed for closed-loop geothermal systems.
@@
Third, due to the limited expressive capacity inherent in fully-connected networks within vanilla PINNs, the training process of vanilla PINNs may lack robustness and stability, potentially hindering convergence to the global minimum. In addition, fully-connected networks struggle to learn multiscale and multiphysics problems. To resolve this issue, more robust NN architectures, such as a multi-scale deep neural network in universal PINNs [104], an ensemble model of PINNs [117], graph neural networks in physics-informed GNNs [118], multi-head physics-informed neural networks (MH-PINNs) [119], a convolutional backbone in Phase2vec [120], and orthogonal polynomials in PI-PINN [121] have been developed. In addition, researchers started to blend physics with RNN-based, CNN-based, or GNN-based learning. For example, a physics-informed convolutional-recurrent learning architecture (PhyCRNet) [116] is proposed for solving PDEs without any labeled data. A physics-constrained CNN learning architecture, PhyGeoNet [122], is developed to learn solutions of parametric PDEs on irregular domains without relying on labeled data. A fast and continuous CNN-based solution (Spline-PINN) [123] is designed for approaching PDE training without any precomputed training data. Graph network [124] is used to represent meshes, and physics-based loss is formulated to provide an unsupervised learning framework for PDEs. However, these methods usually have limited accuracy due to gradient solving at the grid scale (such as finite difference).
@@
Fourth, vanilla PINNs tend to be significantly slower than classical numerical methods. To address this, meta-learning based methods have been proposed for efficiently solving a variety of related PDEs using PINNs. For instance, de Avila Belbute-Peres et al. [125] meta-train a hypernetwork that can generate weights for a small neural network specific to each task, providing an approximate solution to the PDE. Psaros et al. [126] meta-learn a loss function that optimizes the neural network, achieving performance benefits over hand-crafted and online adaptive loss functions. Similarly, Penwarden et al. [127] propose a meta-learning approach similar to our Meta-PDE approach, which learns an initialization of weights to enable quick optimization of the neural network. However, Penwarden et al. [127] find that Model-Agnostic Meta-Learning (MAML) [150] achieves poor performance, only marginally better than random initialization; the recently published Qin et al. [151] come to the opposite conclusion, that MAML-based Meta-PDE performs well for a given runtime. In addition, a cohort of other methods is also proposed to improve the efficiency of PINNs. For instance, XPINNs [128] is presented by space and time parallelization. To replace the computationally expensive automatic differentiation in PINNs, meshless radial basis function-finite differences in DT-PINNs [130] are applied by sparse-matrix vector multiplication, and a coupled-automatic-numerical differentiation framework (canPINN) [129] unifies the advantages of automatic differentiation and numerical differentiation, providing more robust and efficient training than automatic differentiation-based PINNs.
@@
Fifth, in order to improve the generalization of PINNs, a transfer PINN [131] is used to solve forward and inverse problems in nonlinear partial differential equations by parameter sharing. Sixth, PINNs encounter challenges when approximating target functions with high-frequency or multi-scale features. A sea of variants of PINNs are proposed to address this problem. For example, finite basis PINNs [132] use ideas from domain decomposition for complex, multi-scale solutions. Krishnapriyan et al. [133] use curriculum PINN regularization and sequence-to-sequence learning to address the problems of PINNs that fail to capture relevant physical phenomena for slightly more complex problems. In addition, spatio-temporal and multiscale random Fourier features in multiscale PINNs [134] are used to address the high-frequency or multi-scale problems in PINNs. Seventh, adaptive algorithms for weighing components of the PINN loss function (hyperparameter selection) have also been proposed, such as self-adaptation PINNs [135], a heuristic method [136].
@@
Finally, PINNs are employed to address the inverse problem of parameter identification, also known as data-physics-driven parameter discovery. The inverse problem of parameter estimation is often ill-posed due to the non-uniqueness resulting from a large number of unknowns. Moreover, the stability of solutions against data noise and modeling errors is typically not guaranteed. Physics-Informed Neural Networks with Functional Connections (PINN-TFC), such as the extreme theory of functional connections [137], have been developed to mitigate these challenges. PINNs with sparse regression have also been proposed for discovering the governing partial differential equations of nonlinear spatiotemporal systems from sparse and noisy data [152]. This discovery method integrates the strengths of NNs in rich representation learning, physical embedding, automatic differentiation, and sparse regression. Saqlain et al. [138] selected a series of increasingly complex but physically relevant examples and explored the use of PINNs in intrinsically discrete, high-dimensional settings. However, solving such inverse problems using PINNs is still in its early stages, and further research is warranted to advance their applicability.
@@ 2.2.2 Physics-informed Neural Operators
To overcome the challenges of vanilla PINNs, researchers have proposed physics-informed neural operators (PINOs) [140]. PINOs leverage both data and equation constraints, whenever available, within deep operator networks [153]. One notable example is the physics-based DeepONet [139], which trains a DeepONet by integrating known differential equations directly into the loss function along with labeled dataset information. The output of DeepONet is differentiable with respect to its input coordinates, enabling the use of automatic differentiation to develop effective regularization mechanisms that bias the target output function towards satisfying the underlying PDE constraints. However, the PDE loss of physics-informed DeepONet is computed at any query point. The input sensors limit this PINO to a fixed grid or basis, and therefore it is not discretization invariant. In addition, its architecture comes with the limitations of linear approximation. Furthermore, in order to enhance the computational efficiency and generalization of physics-informed DeepONet, a subset of problem-specific surrogate models based on the physics-informed DeepONet is designed, such as physics-informed DeepONet for chemical kinetics [154], and physics-informed variational formulation of DeepONet for brittle fracture analysis [18]. Furthermore, in order to achieve discretization invariance, the physics-informed FNO (PINO) integrates operator learning and physics-informed settings [140]. PINO reduces the need for labeled datasets during neural operator training and facilitates faster convergence of solutions. In this setting, it’s crucial to highlight that, unlike DeepONet, FNO produces the solution on a grid using spectral methods, finite difference, or other numerical gradient solution methods. However, PINO has limited accuracy in solving non-periodic problems due to the numerical gradient method and relies on certain training samples to achieve better accuracy. In addition, the adaptability to different PDE solution problems is not strong due to the selection of model parameters. It cannot be applied to large-scale and long-term sequence PDE solutions. In order to address the problem that PINO cannot represent nonperiodic functions, an architecture that leverages Fourier continuation (FC-PINO) [141] is proposed to apply the exact gradient method to PINO. In addition, a geometry-adaptive physics-informed neural solver [142] is proposed based on the advantages of no training data in physics-informed neural networks; this solver also possesses a fast, accurate, geometry-adaptive and resolution-invariant architecture. Most recently, a physics-informed WNO [143] is proposed for learning the solution operators of families of parametric PDEs without the need for labeled training data, leveraging the advantage of time-frequency localization of wavelets.
@@ 2.3.1 Physical Equation-embedded Machine Learning
PEeML uses the knowledge of specific equations (such as differential equations) to design the frameworks or modules of machine learning. The objectives and limitations of different PEeML are briefly discussed in Table 3.
Specifically, an innovative framework, AutoIP [155], is introduced to automatically integrate physics into Gaussian processes, thereby improving prediction accuracy and uncertainty quantification across diverse differential equations. Additionally, a physics-aware finite volume neural network (FINN) [156] is designed for modeling spatio-temporal advection-diffusion processes by compositively representing the elements of PDEs. Deep Lagrangian networks (DeLaN) [157] encode the Euler-Lagrange equation derived from Lagrangian mechanics. DeLaN can be optimized end-to-end while ensuring adherence to physical principles. A framework called MeshGraphNets [158] is proposed to solve the underlying PDEs using graph neural networks. MeshGraphNets operate by encoding the simulation state into a graph, performing calculations in both the mesh space (defined by the simulation mesh) and the Euclidean world space (where the simulation manifold is embedded). This approach allows the approximation of differential operators essential for the internal dynamics of most physical systems. Furthermore, fluid graph networks (FGN) [159] using graphs to represent the fluid field, is proposed for solving a Lagrangian representation of the Navier-Stokes equation. FGN maintains key physical properties of incompressible fluids, such as low-velocity divergence. Several follow-up works are introduced using graph neural networks [160, 161]. Apart from building an explicit graph structure, an efficient ConvNet architecture based on a continuous convolution layer [44] is developed. The network processes sets of particles in which dynamic particles and static particles are used to represent the fluid and describe the boundary of the scene, respectively.
@@
Table 3: Proposed Physical Equation-embedded Machine Learning
MeshGraphNets [158] Mesh-based partial differential equations A method for learning mesh-based simulations via graph neural networks involves two spaces: the mesh-space (spanned by the simulation mesh) and the Euclidean world-space (where the simulation manifold resides) High computational complexity; hyperparameters
FGN [159] A Lagrangian viewpoint of the Navier-Stokes equation Fluid graph networks (FGN) represent fluid fields using a graph-based approach. In FGN, fluid particles are depicted as nodes, and their interactions are illustrated as edges. The FGN framework decomposes the simulation process into distinct components, such as advection, collision handling, and pressure projection Limited generalization and computational efficiency
Differentiable modules Differential equations These modules implement physical equations as differentiable feature space or computational graphs, enabling the optimization of dynamic processes with analytical gradients and therefore improving sample efficiency. Examples include differentiable physics-informed graph networks [163], a differentiable layer (PDE-Constrained-Layer) [164], physics-aware difference graph networks [165], physics-based finite difference convolution connection [166], and differentiable modeling [13] Limitations on large-scale applications; require an in-depth understanding of differential equations.

### v5

@@ 2.2.1 Physics-informed Neural Networks
PDE solutions are represented as NNs by including the square of the PDE residual in the loss function, resulting in a NN-based PDE solver [24]. Recently, this approach is refined further and called “physics-informed neural networks (PINNs) [25],” initiating a flurry of follow-up work. For example, we consider a parametrized PDE system given by
$\begin{gathered}f\left(\mathbf{x},t,u,\frac{\partial u}{\partial\mathbf{x}},\frac{\partial^{2}u}{\partial\mathbf{x^{2}}},\frac{\partial u}{\partial t},\ldots,\vartheta\right)=0,\quad\mathbf{x}\in\Omega,t\in[0,T],\\ u\left(\mathbf{x},t_{0}\right)=g_{0}(\mathbf{x})\quad\mathbf{x}\in\Omega,\quad u(\mathbf{x},t)=g_{\Gamma}(t)\quad\mathbf{x}\in\partial\Omega,t\in[0,T],\end{gathered}$ (1)
where $u:[0,T]\times\mathbf{X}\rightarrow\mathbb{R}^{n}$ is the solution, subject to initial condition $g_{0}(\mathbf{x})$ and boundary condition $g_{\Gamma}(t)$, which can be of various types such as periodic, Dirichlet, or Neumann. The PDE parameters are denoted by $\vartheta$ = [$\vartheta 1$, $\vartheta 2$, . . . ]. The residual of the PDE is represented by $f$, which includes the differential operators (i.e., $\frac{\partial u}{\partial\mathbf{x}},\frac{\partial^{2}u}{\partial\mathbf{x^{2}}},\frac{\partial u}{\partial t}$, . . . ). The physical domain is denoted by $\Omega$, with its corresponding boundary represented by $\partial\Omega$.
@@
$\begin{gathered}\min_{\theta}\mathcal{L}_{\mathrm{PINN}}(\theta)=w_{i}\mathcal{L}_{IC}+w_{b}\mathcal{L}_{BC}+w_{d}\mathcal{L}_{\text{Data }}+w_{p}\mathcal{L}_{PDE},\\ \mathcal{L}_{IC}=\|u\left(\mathbf{x},t_{0}\right)-g_{0}(\mathbf{x})\|_{\Omega},\quad\mathcal{L}_{BC}=\|u(\mathbf{x},t)-g_{\Gamma}(t)\|_{\partial\Omega},\quad\mathcal{L}_{\text{Data }}=u|_{\Omega}-\hat{u}|_{Data},\\ \mathcal{L}_{PDE}=\|f\left(\mathbf{x},t,u,\frac{\partial u}{\partial\mathbf{x}},\frac{\partial^{2}u}{\partial\mathbf{x^{2}}},\frac{\partial u}{\partial t},\ldots,\vartheta\right)\|_{\Omega},\end{gathered}$ (2)
@@
In current ML-based methods for scientific computing, vanilla PINNs are typically formulated in a pointwise manner, as depicted in Fig. 3(a). These methods utilize NNs as the ansatz for the solution function and optimize a loss function to minimize violation of the given equation, leveraging auto-differentiation for exact, mesh-free derivatives. One of the primary advantages of vanilla PINNs is their ability to avoid the need for a discretization grid, making them particularly useful for solving inverse and higher-dimensional problems. However, despite these advantages, vanilla PINNs exhibit lower efficiency compared to classical methods in solving most PDEs, for several reasons [108, 109]: (1) the challenging optimizations due to soft physics or PDE constraints; (2) the difficulty of conveying information from the initial or boundary conditions to unseen interior regions or to future times; (3) the sensitivity to hyperparameter selection; (4) the slow training/re-optimization process in practical applications and different instances; and (5) failure to converge on time-dependent PDEs or high-frequency, multi-scale PDEs, as shown in Fig. 3(b).
As elucidated in Supplementary Table 2 of SM, a vast and growing body of work aims to address these challenges. Specifically, it encompasses the following eight aspects. First, many problem-specific insights are utilized by improving the training of PINNs. For example, Universal PINNs [110] improve the PINNs by a lower bound constrained uncertainty weighting algorithm and a multi-scale deep NN. A distinct conservative physics-informed neural network (cPINN) [111] is developed for discrete domains dealing with nonlinear conservation laws. Another approach, nPINNs [112], extends PINNs to handle parameter and function inference for integral equations like nonlocal Poisson and nonlocal turbulence models. An improved PINN approach [113] is presented to solve localized wave solutions of the derivative nonlinear Schrödinger equation in complex space, achieving fast convergence and optimal simulation performance. An automatic numerical solver [114] is proposed for the Allen-Cahn and Cahn-Hilliard equations employing advanced PINNs. A simple and efficient characteristic-informed neural network [115] is developed for solving forward and inverse problems in hyperbolic PDEs. NSFnets [107] is proposed for the incompressible Navier-Stokes equations using automatic differentiation. Furthermore, for solving Navier-Stokes equations on irregular geometries, a geometry-aware physics-informed neural network (GAPINN) [116] is presented. In addition, a coupled physics-informed neural networks (CPINNs) [117] is designed for closed-loop geothermal systems.
@@
Third, due to the limited expressive capacity inherent in fully-connected networks within vanilla PINNs, the training process of vanilla PINNs may lack robustness and stability, potentially hindering convergence to the global minimum. In addition, fully-connected networks struggle to learn multiscale and multiphysics problems. To resolve this issue, more robust NN architectures, such as a multi-scale deep neural network in universal PINNs [110], an ensemble model of PINNs [121], graph neural networks in physics-informed GNNs [122], multi-head physics-informed neural networks (MH-PINNs) [123], a convolutional backbone in Phase2vec [124], and orthogonal polynomials in PI-PINN [125] have been developed. In addition, researchers start to blend physics with RNN-based, CNN-based, or GNN-based learning. For example, a physics-informed convolutional-recurrent learning architecture (PhyCRNet) [126] is proposed for solving PDEs without any labeled data. A physics-constrained CNN learning architecture, PhyGeoNet [127], is developed to learn solutions of parametric PDEs on irregular domains without relying on labeled data. A fast and continuous CNN-based solution (Spline-PINN) [128] is designed for approaching PDE training without any precomputed training data. Graph network [129] is used to represent meshes, and physics-based loss is formulated to provide an unsupervised learning framework for PDEs. These models have advanced in exploring network frameworks and effective feature representations. However, they usually have limited accuracy due to gradient solving at the grid scale (such as finite difference).
@@
Fourth, vanilla PINNs tend to be significantly slower than classical numerical methods. To address this, meta-learning based methods have been proposed for efficiently solving a variety of related PDEs using PINNs. For instance, de Avila Belbute-Peres et al. [130] meta-train a hypernetwork that can generate weights for a small NN specific to each task, providing an approximate solution to the PDE. Psaros et al. [131] meta-learn a loss function that optimizes the NN, achieving performance benefits over hand-crafted and online adaptive loss functions. Penwarden et al. [132] propose a meta-learning approach similar to Meta-PDE approach, which learns an initialization of weights to enable quick optimization of the NN. Penwarden et al. [132] find that Model-Agnostic Meta-Learning (MAML) [133] achieves poor performance, only marginally better than random initialization; the recent work [134] comes to the opposite conclusion, that MAML-based Meta-PDE performs well for a given runtime. In addition, a cohort of other methods is also proposed to improve the efficiency of PINNs. For instance, XPINNs [135] is presented by space and time parallelization. To replace the computationally expensive automatic differentiation in PINNs, meshless radial basis function-finite differences in DT-PINNs [136] are applied by sparse-matrix vector multiplication, and a coupled-automatic-numerical differentiation framework (canPINN) [137] unifies the advantages of automatic differentiation and numerical differentiation, providing more robust and efficient training than automatic differentiation-based PINNs.
@@
Fifth, to improve the generalization of PINNs, a transfer PINN [138] is used to solve forward and inverse problems in nonlinear PDEs by parameter sharing. Sixth, PINNs encounter challenges when approximating target functions with high-frequency or multi-scale features. A sea of variants of PINNs are proposed to address this problem. For example, finite basis PINNs [139] use ideas from domain decomposition for complex, multi-scale solutions. Krishnapriyan et al. [140] use curriculum PINN regularization and sequence-to-sequence learning to address the problems of PINNs that fail to capture relevant physical phenomena for slightly more complex problems. In addition, spatio-temporal and multiscale random Fourier features in multiscale PINNs [141] are used to address the high-frequency or multi-scale problems in PINNs. Seventh, adaptive algorithms for weighing components of the PINN loss function (hyperparameter selection) have also been proposed, such as self-adaptation PINNs [142], a heuristic method [143].
@@
Finally, PINNs are employed to address the inverse problem of parameter identification, also known as data-physics-driven parameter discovery. The inverse problem of parameter estimation is often ill-posed due to the non-uniqueness resulting from a large number of unknowns. Moreover, the stability of solutions against data noise and modeling errors is typically not guaranteed. Physics-Informed Neural Networks with Functional Connections (PINN-TFC), such as the extreme theory of functional connections [144], have been developed to mitigate these challenges. PINNs with sparse regression [145] have also been proposed for discovering the governing PDEs of nonlinear spatiotemporal systems from sparse and noisy data. This discovery method integrates the strengths of NNs in rich representation learning, physical embedding, automatic differentiation, and sparse regression. Saqlain et al. [146] select a series of increasingly complex but physically relevant examples and explore the use of PINNs in intrinsically discrete, high-dimensional settings. However, solving such inverse problems using PINNs is still at an early stage, and further studies are certainly warranted.
@@ 2.2.2 Physics-informed Neural Operators
To overcome the challenges of vanilla PINNs, physics-informed neural operators (PINOs) [147] are proposed. These utilize both the data and equation constraints (whichever are available) for deep operator networks [148].
One particular PINO, physics-informed DeepONet [149], trains the DeepONet by integrating known differential equations directly into the loss function along with labeled dataset information. The output of DeepONet is differentiable with respect to its input coordinates, enabling the use of automatic differentiation to develop effective regularization mechanisms that bias the target output function towards satisfying the underlying PDE constraints. However, the PDE loss of physics-informed DeepONet is computed at any query point. The input sensors limit this PINO to a fixed grid or basis, and therefore it is not discretization invariant. In addition, its architecture comes with the limitations of linear approximation. Furthermore, in order to enhance the computational efficiency and generalization of physics-informed DeepONet, a subset of problem-specific surrogate models based on the physics-informed DeepONet is designed, such as physics-informed DeepONet for chemical kinetics [150], and physics-informed variational formulation of DeepONet for brittle fracture analysis [18]. Furthermore, in order to achieve discretization invariance, the physics-informed FNO (PINO) [147] is proposed by integrating operator learning and physics-informed settings. PINO reduces the need for labeled datasets during neural operator training and facilitates faster convergence of solutions. In this setting, it’s crucial to highlight that, unlike DeepONet, FNO produces the solution on a grid using spectral methods, finite difference, or other numerical gradient solution methods. However, PINO has limited accuracy in solving non-periodic problems due to the numerical gradient method and relies on certain training samples to achieve better accuracy. In addition, the adaptability to different PDE solution problems is not strong due to the selection of model parameters. It cannot be applied to large-scale and long-term sequence PDE solutions. In order to address the problem that PINO cannot represent nonperiodic functions, an architecture that leverages Fourier continuation (FC-PINO) [151] is proposed to apply the exact gradient method to PINO. In addition, a geometry-adaptive physics-informed neural solver (GeoPINS) [35] is proposed based on the advantages of no training data in physics-informed neural networks; this solver also possesses a fast, accurate, geometry-adaptive and resolution-invariant architecture. Most recently, a physics-informed WNO [152] is proposed for learning the solution operators of families of parametric PDEs without the need for labeled training data, leveraging the advantage of time-frequency localization of wavelets.
@@ 2.3.1 Physical Equation-embedded Machine Learning
PEeML uses the knowledge of specific equations (such as differential equations) to design the frameworks or modules of ML. The objectives and limitations of different PEeML are briefly discussed in Supplementary Table 3 of SM.
Specifically, a powerful framework, AutoIP [153], is introduced to automatically integrate physics into Gaussian processes, thereby improving prediction accuracy and uncertainty quantification across diverse differential equations. Additionally, a physics-aware finite volume neural network (FINN) [154] is designed for modeling spatio-temporal advection-diffusion processes by compositively representing the elements of PDEs. Deep Lagrangian networks (DeLaN) [155] encode the Euler-Lagrange equation derived from Lagrangian mechanics. DeLaN can be optimized end-to-end while ensuring adherence to physical principles. A framework called MeshGraphNets [156] is proposed to solve the underlying PDEs using graph neural networks. MeshGraphNets works by encoding the simulation state into a graph, performing calculations in both the mesh space (defined by the simulation mesh) and the Euclidean world space (where the simulation manifold is embedded). This approach allows the approximation of differential operators essential for the internal dynamics of most physical systems. Furthermore, fluid graph networks (FGN) [157] using graphs to represent the fluid field, is proposed for solving a Lagrangian representation of the Navier-Stokes equation. FGN maintains key physical properties of incompressible fluids, such as low-velocity divergence. Several follow-up works are introduced using graph neural networks [158, 159]. Apart from building an explicit graph structure, an efficient ConvNet architecture based on a continuous convolution layer [76] is developed. The network processes sets of particles in which dynamic particles and static particles are used to represent the fluid and describe the boundary of the scene, respectively.